OpenStudy (anonymous):

The sum of the interior angles of a polygon is the same as the sum of its exterior angles. What type of polygon is it?

OpenStudy (tkhunny):

Remember that interior and exterior angles are supplementary. This solves the problem.

OpenStudy (merchandize):

360 = (n-2)(180) n = 2+2 = 4 quadrilateral